OpenStudy (haleyelizabeth2017):

Suppose the second hand on a clock is 3 inches long. Find the linear velocity of the tip of the second hand.

jimthompson5910 (jim_thompson5910):

what is the circumference of a circle with radius 3?

OpenStudy (haleyelizabeth2017):

6pi or 18.8

jimthompson5910 (jim_thompson5910):

let's stick with 6pi since it's exact

OpenStudy (haleyelizabeth2017):

okay

jimthompson5910 (jim_thompson5910):

by definition, the second hand moves around the full circle in 60 seconds (assuming the clock is working properly) so it covers 6pi inches in 60 seconds we can form this fraction \[\Large \frac{6 \pi \ \text{ inches}}{60 \ \text{ seconds}}\] to represent the speed of the tip of the second hand

OpenStudy (haleyelizabeth2017):

Ohh!

jimthompson5910 (jim_thompson5910):

now simply reduce that fraction \[\Large \frac{6 \pi \ \text{ inches}}{60 \ \text{ seconds}} = \frac{\pi \ \text{ inches}}{10 \ \text{ seconds}}\]

jimthompson5910 (jim_thompson5910):

the exact linear speed is pi/10 inches per second approx that is 0.31415926535898 inches per second since pi/10 = 0.31415926535898

OpenStudy (haleyelizabeth2017):

Thank you so much, I was so confused at first when I read this :)

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (haleyelizabeth2017):

:)